OpenStudy (anonymous):

Help Please!Create a graph for the following situation. Jessica has several horses that she needs to feed daily. Each day she uses 2 bales of hay. After 7 days she has 1 hay bale remaining. Be sure to identify the slope and y-intercept.

OpenStudy (anonymous):

So, you have a linear graph with slope of -2. She looses 2 bales of hay ever day, and has one at the end. so x - 2(7) = 1 x -14 = 1 x = 15 So, your y intercept is 15, and your slope is -2. can you graph that?

OpenStudy (anonymous):

OMG! I'm going to try right now.

OpenStudy (anonymous):

I think we still have a problem... I don't think my son is graphing it right

OpenStudy (anonymous):

If we use trogdortheburninator's numbers we have y = -2x + 15. Okay so on a graph you would be starting at 0,15. Then for every number you would go down two and over one, so the next point would be 1, 13. You'll have to do this another 6 times I think if I understand the question correctly.

OpenStudy (anonymous):

The graph that was provided by the teacher does not allow for him to go down...

OpenStudy (anonymous):

is it a multiple choice question?

OpenStudy (anonymous):

This is the question; Create a graph for the following situation. Jessica has several horses that she needs to feed daily. Each day she uses 2 bales of hay. After 7 days she has 1 hay bale remaining. Be sure to identify the slope and y-intercept.

OpenStudy (anonymous):

Hmm..I know that it has to go into y = mx + b form..and I'm almost positive that the 2 is the slope, but I can't seem to get it to work out -- sorry. @Mertsj could you please help this person?

OpenStudy (anonymous):

I appreciate your help anyway! Thanks :)

OpenStudy (mertsj):

The slope is NEGATIVE 2, not 2. So the equation is y = -2x+15 The y intercept is 15 since at day 0 she had 15 bales of hay.

OpenStudy (mertsj):